OpenStudy (anonymous):

3x + y = 5 2x – y = 10

OpenStudy (anonymous):

?

OpenStudy (anonymous):

are you trying to solve it?

OpenStudy (anonymous):

@no12 what's the question? :/

OpenStudy (jhannybean):

I believe there is a plus instead of an equal sign infront of that "52x"

OpenStudy (anonymous):

And i guess she wants to solve it with the help of any method like elimination......

OpenStudy (anonymous):

No, I think he just doesn't check to see how his post got formatted. If you copy it to a text editor, you'll find a space between the 2 and the 5. I think he means: 3x + y = 5 2x - y = 10

OpenStudy (anonymous):

Which he can solve by simply adding them. :-)

OpenStudy (anonymous):

No no, the text editor wanted to write like this only but wanted to solve by any method not by adding :/

OpenStudy (anonymous):

But adding them eliminates the y, leaving the x value. After finding x, he can then find y by substituting x

OpenStudy (anonymous):

2x-y=10 3x+y=5 5x=15 x=3 3(3)+y=5 9+y=5 y=-4 (3,-4)

OpenStudy (anonymous):

Yes. Like that. :-)